# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

### v1

@@ 2.1 Motion Compensation-based
Before the advent of deep-learning, VFI was primarily tackled using Motion-Compensated Frame Interpolation (MCFI) [64, 67, 68, 70] or Frame Rate Up-Conversion (FRUC) [63, 65, 66, 69, 71, 72], which dominated the field from the late 1990s through the early 2000s. These approaches explicitly estimate motion, typically via block matching or global parametric models, and synthesize the intermediate frame by warping the input frames according to the estimated motion fields. A typical MCFI pipeline involves two steps: (i) block-based motion estimation and (ii) pixel-level warping for frame synthesis. In block-based estimation, the frame is partitioned into fixed-size rectangular blocks, assuming uniform motion within each region. While this formulation offers computational efficiency, it fails to capture non-rigid or object-specific motion, often resulting in artifacts such as holes (due to occlusions) and overlaps (due to many-to-one mappings). Rooted in classical video coding frameworks [121], MCFI methods emphasize speed and simplcity, but inherently lack the capacity to handle fine-grained, non-linear motion. To address these issues, several extensions are proposed, including multi-stage estimation [70], adaptive motion models [66], and occlusion-aware warping [72]. Intermediate frame synthesis is typically achieved via block-wise projection or forward guided by the estimated motion.
@@
Despite their limited robustness under complex dynamics, MCFI and FRUC methods [63, 64, 65, 66, 67, 68, 69, 70, 71, 72] lay the conceptual foundation for modern VFI. This core principle, which involves explicit motion estimation followed by motion-guided warping, remains central to many modern learning-based models and is now enhanced with deep feature representations and end-to-end training. Importantly, classical motion-compensated strategies offer valuable insights into the inductive biases that shape modern VFI architectures. Concepts such as motion locality, piecewise rigidity, and spatial warping, which originated from block-based estimation, are implicitly retained in modern mechanisms including deformable convolutions [122, 32] and local attention [123]. Furthermore, the challenges encountered in this approach, such as occlusion handling and motion discontinuity, have directly influenced the design of occlusion-aware blending modules and bidirectional flow formulations in recent models. In this light, traditional motion models serve as both a historical foundation and a conceptual framework for the progressive development of VFI architectures.
2.2 Deep Learning-based
2.2.1 Kernel-based
@@
Kernel-based VFI methods [28, 29, 16, 30, 31, 32, 33, 34, 35, 36, 37, 8, 38, 39, 40, 41, 42, 43, 44, 45, 46, 56, 47] synthesize intermediate frames by predicting spatially-adaptive convolutional kernels, which are subsequently applied to local patches extracted from the input frames. Motion information is implicitly encoded in the kernel weights, thereby enabling motion-aware pixel aggregation without explicit motion estimation. As shown in Fig. 5 (a), a standard kernel-based interpolation can be mathematically formulated as:
where $N$ denotes the number of input frames, $R$ denotes the kernel size, and $W_{k,l}$ represents the learned kernel weight at offset $(k,l)$. This approach adopts a simple single-stage formulation that combines motion estimation and frame synthesis into a one-step process. AdaConv [28] utilizes a U-Net-like architecture [24] to predict spatially-varying 2D kernels for each output pixel. This enables local, pixel-wise motion-aware aggregation that can implicitly handle both alignment and occlusion [46]. SepConv [29] further reduces the computational overhead by decomposing the 2D kernel into separable 1D kernels:
$W=W_{v}*W_{h},$ (3)
where $W_{v}\in\mathbb{R}^{R\times 1}$ and $W_{h}\in\mathbb{R}^{1\times R}$ are vertical and horizontal 1D kernels respectively. The * denotes the outer product between the two 1D kernels, resulting in a full 2D kernel $W\in\mathbb{R}^{R\times R}$. This decomposition reduces the number of learnable parameters from $R^{2}$ to $2R$, while maintaining a comparable receptive field. Despite their conceptual simplicity, these methods are inherently limited in handling large displacements due to their fixed receptive fields [74]. Such constraints stem from the content-agnostic nature of CNNs, which uniformly apply learned filters across spatial locations [51]. While this weight-sharing inductive bias proves effective in recognition tasks, it becomes suboptimal in VFI, where fine-grained motion modeling is essential. To overcome this problem, deformable kernel-based methods [32, 34, 37, 8, 40, 42, 39, 44, 51, 56] introduce learnable offsets [122] as shown Fig. 5 (b), which allow sampling outside the regular convolution grid:
$\hat{I}_{(}x,y)=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}$
$\cdot I_{i}(x+k+\alpha_{k,l},\ y+l+\beta_{k,l}),$ (4)
where $(\alpha_{k,l},\beta_{k,l})$ are the learnable offsets. AdaCoF [36] estimates both kernel weights and sampling offsets for each output pixel, though it employs a fixed offset pattern, limiting its expressiveness under complex motion. To enhance spatial adaptivity further as shown in Fig. 5 (c), dynamic kernel-based methods [58, 35, 41, 44, 124] make location-dependent kernel weights:
$\hat{I}(x,y)=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}(x,y)$
$\cdot I_{i}(x+k+\alpha_{k,l},\ y+l+\beta_{k,l}),$ (5)
where $W_{k,l}(x,y)$ denotes a dynamically predicted kernel at location $(x,y)$. Methods such as CDFI [41] and MSEConv [44] jointly learn spatially-varying weights and offsets, offering enhanced flexibility and improved interpolation accuracy.
Overall, kernel-based methods forego explicit motion supervision, instead leveraging learned spatial priors for synthesizing intermediate frames. As kernel prediction and convolution-based synthesis are tightly coupled, motion estimation and frame synthesis are implicitly fused into a single-stage. This implicit formulation offers robustness in noisy or uncertain motion settings and eliminates the dependency on accurate optical flow. However, these models often lack temporal generalizability, as the learned kernels are typically conditioned on a fixed interpolation time (e.g., $t=0.5$). Consequently, most kernel-based methods are constrained to CTFI (Sec. 3.1) and fail to generalize to ATFI (Sec. 3.2), limiting their applicability in real-world scenarios requiring temporal flexibility.
2.2.2 Flow-based
@@
Flow-based methods [16, 73, 74, 75, 76, 62, 60, 77, 78, 79, 39, 80, 81, 58, 59, 82, 83, 84, 125, 55, 85, 54, 53, 87, 88, 89, 90, 30, 33, 35, 37, 38, 47] explicitly estimate dense motion in the form of optical flow, a dense motion field representing the pixel-wise displacements between two frames, to temporally align input frames and synthesize intermediate frames. Recent advances in optical flow estimation [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] have directly propelled the performance of flow-based VFI models. A typical pipeline comprises: (1) estimating either anchor flows ($\mathcal{V}_{0\rightarrow t},\mathcal{V}_{1\rightarrow t}$) or intermediate flows ($\mathcal{V}_{t\rightarrow 0},\mathcal{V}_{t\rightarrow 1}$), (2) applying flow-guided warping [60, 61] of input frames or features ($I_{0},I_{1}$ or $F_{0},F_{1}$), and (3) synthesizing the target frame ($\hat{I}_{t}$) by blending the warped results ($\hat{I}_{0\rightarrow t}/\hat{F}_{0\rightarrow t}$ and $\hat{I}_{1\rightarrow t}/\hat{F}_{1\rightarrow t}$).
@@
The accuracy of the flow critically impacts interpolation quality in this approach, as misalignment directly causes blur and artifacts. Many earlier works [30, 7, 60] adopt off-the-shelf optical flow networks [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] to estimate the initial flows. Although these networks offer reliable motion estimation, they are not specifically optimized for the VFI task and often introduce unnecessary architectural complexity. Moreover, they tend to have large model sizes and struggle to handle extreme motions that lie outside the training distribution [125]. To better adapt the motion estimation to the VFI task, a number of methods [16, 5, 6, 74, 62, 35, 59, 81, 97, 83, 55, 54, 53, 125, 57, 100] propose to estimate their own task-oriented flow within their framework, which is optimized jointly with the frame interpolation objective. For example, BiM-VFI [137] distills flow knowledge from an ensemble of flow predictors into a lightweight network tailored for interpolation. GIMM-VFI [88] addresses the noise in flows from pre-trained optical flow estimator (e.g., RAFT [134], FlowFormer [135]) by refining them through a coordinate-based implicit networks. Pseudo ground-truth (GT) strategies are also common, where pseudo GT flow is generated by existing flow networks and used as weak supervision to bootstrap VFI training [16, 79]. These help produce temporally consistent and semantically aligned flows customized for interpolation. With the estimated flow, warping is implemented via either forward [60] or backward [61] warping operation. Depending on the warping operation, it requires different types of flows.
@@
Figure 6: Comparison of forward and backward warping strategies. (a) Forward warping [60] projects source pixels ($I_{0}$) to their estimated positions in the target frame using $\mathcal{V}_{0\rightarrow t}$. This may introduce holes (unmapped pixels) or collisions (multiple pixels mapped to the same location). (b) Backward warping [61] samples each pixel in the target frame from the source using $\mathcal{V}_{t\rightarrow 0}$. Since sampling is performed at every target location, backward warping naturally produces dense and complete outputs.
Forward warping. Most forward warping-based methods [30, 6, 73, 74, 62, 60, 35, 38, 80, 125, 85] first estimate bidirectional flows $(\mathcal{V}_{0\rightarrow 1},\mathcal{V}_{1\rightarrow 0})$, from which the intermediate flows $(\mathcal{V}_{0\rightarrow t},\mathcal{V}_{1\rightarrow t})$ are linearly interpolated:
@@
These flows are used to project source pixels to the intermediate frame:
@@
However, forward warping introduces structural artifacts such as holes (unmapped regions) and collisions (multiple pixels mapping to the same target position) particularly near motion boundaries [5] as shown in Fig. 6 (a). Unlike backward warping, which ensures dense sampling by mapping each pixel in the target domain, forward warping does not guarantee full coverage due to its source-driven formulation. This intrinsic asymmetry stems from the lack of inverse consistency in optical flow, i.e., $\mathcal{V}_{i\rightarrow j}\neq-\mathcal{V}_{j\rightarrow i}$ in general, particularly under occlusions or non-rigid motion. To address this, SoftSplat [60] proposes a softmax-based splatting mechanism:
@@
where $Z$ denotes a learned importance map (e.g., depth), and the operator $\vec{\sum}$ denotes a differentiable splatting with soft aggregation. The soft aggregation scheme in SoftSplat not only mitigates hole/collision artifacts but also improves the gradient flow by making warping fully differentiable, in contrast to standard splatting operations which are piecewise constant and non-smooth. Despite this, the inherent artifacts make naive forward warping a less favored primary choice.
Backward warping. In contrast, backward warping [5, 33, 76, 81, 59, 55] samples each pixel in the target frame by mapping it back to the input using estimated intermediate flows $(\mathcal{V}_{t\rightarrow 0},\mathcal{V}_{t\rightarrow 1})$. Intermediate flows denote the flows from the unknown target frame to the input frames. Since the target frame is unavailable, it is not straightforward to obtain the intermediate flows. These flows can be approximated via direct prediction [16, 6, 35, 58, 59, 52, 81, 54, 53, 55, 57, 100, 90], flow interpolation [5, 7], or flow reversal [76, 77, 79, 89]. For instance, SuperSloMo [5] employs such linear approximations and further refines them via dedicated subnetworks. The flow interpolation for intermediate flows is defined as:
$\hat{\cal V}_{t\rightarrow 0}$ $=-t\cdot{\cal V}_{0\rightarrow 1}\,\,\textrm{ or }\,\,t\cdot{\cal V}_{1\rightarrow 0}$ (9)
$\hat{\cal V}_{t\rightarrow 1}$ $=(1-t)\cdot{\cal V}_{0\rightarrow 1}\,\,\textrm{ or }\,\,-(1-t)\cdot{\cal V}_{1\rightarrow 0}.$ (10)
To enhance robustness, XVFI [79] introduces Complementary Flow Reversal (CFR), a weighted aggregation strategy that fuses multiple reversed and complementary flows to construct robust intermediate motion fields. This strategy complements the shortcomings of both linear flow approximation and naive flow reversal [76], offering robustness against ambiguities near motion boundaries. Given the intermediate flows, backward warping is applied as:
@@
where $\overset{\leftarrow}{\mathcal{W}_{b}}$ denotes the backward warping operator [61]. The warped results are blended using occlusion-aware mask $M$ and residual refinement $R$:
$I_{t}=M\odot\hat{I}_{0\rightarrow t}+(1-M)\odot\hat{I}_{1\rightarrow t}+R,$ (12)
@@
The operator $\odot$ denotes element-wise multiplication, or the Hadamard product, which blends the warped frames proportionally based on the occlusion-aware confidence map. Additionally, some methods [5, 76, 75, 79] further incorporates $(1{-}t)$ and $t$ as scalar weights into $M$ to guide time-aware blending. Several methods also exploit auxiliary priors such as depth [7], contextual features [30, 33, 7, 35, 60, 39], or edge information [73, 77, 37] to further guide interpolation. Also, learnable synthesis networks [138] or additional post-processing can further improve sharpness and correct residual artifacts.
@@
Modeling non-linear motion. Many early methods assume linear motion and brightness constancy [16, 5, 30, 33, 7, 73, 76, 74, 6, 35, 60, 85], meaning that objects move along a straight trajectories at constant speed, and pixel intensities remain unchanged. However, these assumptions often fail under real-world scenarios. Quadratic [76, 77, 139, 86] or cubic [62] motion modeling has been proposed to account for acceleration. QVI [76] and EQVI [77] estimate acceleration-aware flows utilizing four input frames. While recent works [137, 90] further explore velocity ambiguity [90], which refers to the ill-posed nature of intermediate motion inference where multiple trajectories yield the same intermediate position, especially under occlusion or acceleration. BimVFI [137] and Zhong et al [90] introduce bidirectional motion fields and time-aware reasoning mechanisms to disambiguate such cases, enabling robust interpolation under occlusion, acceleration, and non-linear motion.
@@
Overall, flow-based methods remain one of the most extensively explored and practically adopted approaches in VFI, owing to their explicit and interpretable modeling of motion trajectories. Their ability to flexibly generate intermediate frames for arbitrary timestamps makes them well-suited for various real-world scenarios requiring variable frame-rate synthesis. Despite these strengths, their performance is sensitive to flow estimation accuracy, particularly under conditions of occlusion, large motion, lighting variation or non-linear motion. As research in optical flow continues to evolve [140, 141], flow-based VFI is expected to further benefit from these developments and remain a foundational component of future VFI approach.
2.2.3 Kernel- and Flow-based Combined

### v3

@@ II-A Motion Compensation-based
A typical MCFI pipeline involves two key steps: (i) block-based motion estimation and (ii) pixel-level warping for frame synthesis. In block-based estimation, each frame is partitioned into fixed-size rectangular blocks under the assumption of uniform motion within each block. While computationally efficient, this design often fails to capture non-rigid or object-specific motion, often resulting in artifacts such as holes (due to occlusions) and overlaps (due to many-to-one mappings). Rooted in classical video coding frameworks [101], MCFI methods emphasize speed and simplicity, but inherently lack the capacity to handle fine-grained, non-linear motion. To address these issues, various extensions have been proposed, including multi-stage motion estimation [90], adaptive motion models [125], and occlusion-aware warping [229]. Intermediate frame synthesis was generally performed through block-wise projection or forward warping, using the estimated motion vectors to guide the placement of each pixel. However, these methods typically operate in the pixel or block space without modeling complex motion patterns, and thus struggle to maintain spatial consistency under non-linear dynamics.
@@
Despite their limited robustness in handling complex dynamics, MCFI and FRUC methods [76, 164, 20, 125, 75, 35, 114, 90, 230, 229] laid the conceptual foundation for modern VFI. Their core principle, explicit motion estimation followed by motion-compensated warping, remains central to many modern learning-based models and is now enhanced with deep feature representations and end-to-end training. Importantly, classical motion-compensated strategies introduced valuable insights into the inductive biases that shape modern VFI architectures. Concepts such as motion locality, piecewise rigidity, and spatial warping, which originated from block-based estimation, are implicitly retained in modern mechanisms like deformable convolutions [40, 288] and local attention [225]. Furthermore, challenges faced by early approaches, such as occlusion handling and motion discontinuity, have directly motivated the development of occlusion-aware blending, bidirectional flow formulations in contemporary VFI models. In this light, traditional motion models serve as both a historical foundation and conceptual framework for the progressive development of VFI architectures.
II-B Deep Learning-based
II-B1 Kernel-based
Kernel-based VFI methods [169, 170, 167, 182, 288, 7, 32, 177, 124, 68, 244, 171, 202, 33, 52, 31, 44, 53, 113, 283, 266, 199] synthesize intermediate frames by predicting spatially-adaptive convolutional kernels, which are applied to local patches extracted from the input frames. Motion information is implicitly encoded in these kernel weights, enabling motion-aware pixel aggregation. In other words, this approach still estimates and exploits motion, but represents it implicitly through the spatial pattern and support of the learned kernels rather than as an explicit dense flow field or motion-vector map. A standard kernel-based interpolation can be formulated as
where $N$ is the number of input frames, $R$ is the kernel size, and $W_{k,l}$ denotes the learned kernel weight at offset $(k,l)$, as shown in Fig. 5 (a). AdaConv [169] utilizes a U-Net-like architecture [191] to predict spatially-varying 2D kernels for each output pixel. This enables local, pixel-wise motion-aware aggregation that can implicitly handle both alignment and occlusion [283]. SepConv [170] further reduces the computational overhead by decomposing the 2D kernel into separable 1D kernels:
$W=W_{v}*W_{h},$ (3)
where $W_{v}\in\mathbb{R}^{R\times 1}$ and $W_{h}\in\mathbb{R}^{1\times R}$ are vertical and horizontal 1D kernels respectively. The * denotes the outer product between the two 1D kernels, resulting in a full 2D kernel $W\in\mathbb{R}^{R\times R}$. This reduces the number of learnable parameters from $R^{2}$ to $2R$, while preserving a comparable receptive field. Despite their simplicity, these methods are inherently limited in handling large displacements due to their fixed receptive fields [262]. Such constraints stem from the content-agnostic nature of CNNs, which uniformly apply learned filters across spatial locations [203]. While this weight-sharing inductive bias proves effective in recognition tasks, it becomes suboptimal in VFI, where fine-grained motion modeling is essential. To overcome this problem, deformable kernel-based methods [288, 32, 68, 244, 33, 31, 202, 53, 203, 266] introduce learnable offsets [40] as shown Fig. 5 (b), which allow sampling beyond the static grid:
$\hat{I}(x,y)=\$$\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}$
$\cdot I_{i}(x+k+\alpha_{k,l},\ y+l+\beta_{k,l}),$ (4)
where $(\alpha_{k,l},\beta_{k,l})$ are the learnable offsets. From a motion-representation viewpoint, these offsets can be interpreted as kernel-centered displacement vectors, while the associated kernel weights modulate how much each sampled location contributes, yielding an implicit analogue of a local flow field. AdaCoF [124] jointly predicts both kernel weights and sampling offsets for each output pixel, improving flexibility over static kernels. However, its limited offset range and time-invariant sampling pattern limits its expressiveness under complex motion. To further enhance spatial adaptivity, dynamic kernel-based methods [178, 177, 52, 53, 138] as shown in Fig. 5 (c) generate location-dependent kernel weights:
$\hat{I}(x,y)=\$$\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}(x,y)$
$\cdot I_{i}(x+k+\alpha_{k,l},\ y+l+\beta_{k,l}),$ (5)
where $W_{k,l}(x,y)$ denotes a dynamically predicted kernel at location $(x,y)$. Methods such as CDFI [52] and MSEConv [53] jointly learn spatially-varying weights and offsets, resulting in enhanced flexibility and improved interpolation accuracy.
@@ II-B1 Kernel-based
Kernel-based models adopt a simple single-stage formulation that combines motion estimation and frame synthesis into a one-step process [169, 170]. Instead of first regressing an explicit flow field and then warping the inputs, the network directly predicts sampling kernels whose spatial support implicitly specifies where information is gathered from the input frames, thereby coupling motion inference and reconstruction in a unified operation. This implicit formulation enhances robustness in motion-ambiguous or low-texture regions by avoiding reliance on external optical flow estimators. However, a notable limitation arises from their temporal rigidity. Most methods are trained to interpolate at fixed time steps (e.g., $t=0.5$) and lack generalization to arbitrary times $t\in(0,1)$. As a result, they are typically restricted to CTFI (Sec. III-A) and fail to support ATFI (Sec. III-B), limiting their applicability in real-world scenarios requiring temporal flexibility.
II-B2 Flow-based
@@
Flow-based methods [179, 148, 145, 262, 187, 251, 34, 168, 144, 271, 204, 202, 89, 120, 178, 93, 196, 188, 166, 111, 176, 109, 133, 270, 104, 73, 240, 281, 167, 7, 177, 68, 171, 199, 17] explicitly estimate dense motion in the form of optical flow, a dense motion field representing the pixel-wise displacements between two frames, and use it to align inputs temporally for intermediate frame synthesis. Advances in optical flow estimation [56, 99, 212, 218] have directly propelled the performance of flow-based VFI models. A typical pipeline consists of three stages: (1) estimating either anchor flows ($\mathcal{V}_{0\rightarrow t},\mathcal{V}_{1\rightarrow t}$) or intermediate flows ($\mathcal{V}_{t\rightarrow 0},\mathcal{V}_{t\rightarrow 1}$), (2) warping [168, 100] the inputs or their features ($I_{0},I_{1}$ or $F_{0},F_{1}$) using the predicted flow fields, and (3) synthesizing the target frame ($\hat{I}_{t}$) by blending the warped results ($\hat{I}_{0\rightarrow t}/\hat{F}_{0\rightarrow t}$ and $\hat{I}_{1\rightarrow t}/\hat{F}_{1\rightarrow t}$).
@@
The accuracy of the flow critically impacts interpolation quality in this approach, as misalignment directly causes blur and artifacts. Early works [167, 6, 168] adopt off-the-shelf optical flow networks [108, 236, 56, 99, 186, 212, 96, 8, 218, 91, 250] to estimate the initial flows. While these networks offer strong general-purpose motion estimation, they are not specifically optimized for the VFI task and struggle to handle motions that lie outside the training distribution [111]. Moreover, their large parameter count introduces unnecessary overhead. To address these issues, a number of methods [148, 107, 252, 262, 34, 177, 93, 120, 46, 188, 176, 133, 270, 111, 139, 267] estimate their own task-oriented flow within their framework, which is optimized jointly with the frame interpolation objective. BiM-VFI [194] distills flow knowledge from an ensemble of flow networks into a compact flow estimator aligned with the interpolation task. GIMM-VFI [73] mitigates the noise in flows from pre-trained flow estimator (e.g., RAFT [218], FlowFormer [91]) by refining them through a coordinate-based implicit networks. Pseudo ground-truth (GT) strategies are also common, where pseudo GT flow is generated by existing flow networks and used as weak supervision to bootstrap VFI training [148, 204]. These help produce temporally consistent and semantically aligned flows customized for interpolation. With the estimated flow, warping is implemented via either forward [168] or backward [100] warping. Depending on the warping operation, it requires different types of flows.
@@
Figure 6: Comparison of forward and backward warping strategies. (a) Forward warping [168] projects source pixels ($I_{0}$) to their estimated positions in the target frame using $\mathcal{V}_{0\rightarrow t}$. (b) Backward warping [100] samples each pixel in the target frame from the source using $\mathcal{V}_{t\rightarrow 0}$.
Forward warping. Forward warping-based methods [167, 252, 145, 262, 34, 168, 177, 171, 89, 111, 109] first estimate bidirectional flows $(\mathcal{V}_{0\rightarrow 1},\mathcal{V}_{1\rightarrow 0})$, from which the intermediate flows $(\mathcal{V}_{0\rightarrow t},\mathcal{V}_{1\rightarrow t})$ are linearly interpolated:
@@
These flows are then used to project source pixels to the target frame:
@@
While conceptually simple, forward warping suffers from two inherent issues: holes (unmapped regions) and collisions (multiple pixels mapped to the same location), especially near motion boundaries [107], as shown in Fig. 6 (a). These artifacts arise from the intrinsic asymmetry of optical flow, where inverse consistency does not generally hold (i.e., $\mathcal{V}_{i\rightarrow j}\neq-\mathcal{V}_{j\rightarrow i}$), particularly under occlusions or non-rigid motion. Moreover, forward warping is generally non-differentiable due to discrete sampling, hindering gradient-based optimization. To address this, SoftSplat [168] proposes a differentiable softmax-based splatting mechanism:
@@
where $Z$ denotes a learned importance map (e.g., depth), and $\vec{\sum}$ denotes a differentiable splatting with soft aggregation. This formulation not only mitigates the aforementioned artifacts but also improves the gradient flow by making warping fully differentiable, in contrast to standard splatting operations which are piecewise constant and non-smooth. Despite this, the inherent artifacts make naive forward warping a less favored primary choice.
Backward warping. Backward warping [107, 7, 251, 120, 93, 176] reconstructs each target pixel by sampling from the input frames using estimated intermediate flows $(\mathcal{V}_{t\rightarrow 0},\mathcal{V}_{t\rightarrow 1})$, which represent motion from the unknown target frame to each input frame. Since the target frame is unavailable, it is not straightforward to obtain these flows. It can be approximated via direct prediction [148, 252, 177, 178, 93, 150, 120, 133, 270, 176, 139, 267, 281], flow interpolation [107, 6], or flow reversal techniques [251, 144, 204, 240]. For instance, SuperSloMo [107] employs linear approximations and further refines them via dedicated subnetworks:
[…]
To enhance robustness against ambiguities near motion boundaries, XVFI [204] introduces Complementary Flow Reversal (CFR), which aggregates multiple reversed flows to construct more stable motion fields. Given the intermediate flows, backward warping is applied as:
@@
where $\overset{\leftarrow}{\mathcal{W}_{b}}$ denotes the backward warping operator [100]. The warped results are blended using occlusion-aware mask $M$ and residual refinement term $R$:
$I_{t}=M\odot\hat{I}_{0\rightarrow t}+(1-M)\odot\hat{I}_{1\rightarrow t}+R.$ (12)
@@
The operator $\odot$ denotes element-wise multiplication, or the Hadamard product, which blends the warped frames proportionally based on the occlusion-aware confidence map. Some methods [107, 251, 187, 204] further incorporates (1t)(1{-}t) and $t$ as scalar weights into $M$ to guide time-aware blending. Several methods also exploit auxiliary priors such as depth [6], contextual features [167, 7, 6, 177, 168, 202], or edge information [145, 144, 68] to further guide interpolation. Learnable synthesis networks [61] are also commonly employed to further sharpen the output and correct residual artifacts.
@@
Modeling non-linear motion. Many early methods [148, 107, 167, 7, 6, 145, 262, 252, 177, 168, 109] assume linear motion and brightness constancy, meaning that objects move along a straight trajectories at constant speed, and pixel intensities remain unchanged. However, these assumptions often fail under real-world scenarios involving acceleration, occlusion, or dynamic lighting. Quadratic [251, 144, 276, 88] or cubic [34] motion modeling has been proposed to account for acceleration. QVI [251] and EQVI [144] estimate acceleration-aware flows utilizing four input frames. While recent works [194, 281] further explore velocity ambiguity [281], which refers to the ill-posed nature of intermediate motion inference where multiple trajectories yield the same intermediate position, especially under occlusion or acceleration. BiM-VFI [194] and Zhong et al [281] introduce bidirectional motion fields and time-aware reasoning mechanisms to disambiguate such cases, enabling robust interpolation under occlusion, acceleration, and non-linear motion.
@@
Overall, flow-based methods remain one of the most extensively explored and practically adopted approaches in VFI, owing to their explicit and interpretable modeling of motion trajectories. Their ability to flexibly generate intermediate frames for arbitrary timestamps makes them well-suited for applications such as variable frame-rate generation and slow-motion rendering. Despite these strengths, their performance is sensitive to flow estimation accuracy, particularly under conditions of occlusion, large motion, lighting variation or non-linear motion. As research in optical flow continues to evolve [57, 55, 10], flow-based VFI is expected to further benefit from these developments and remain a foundational component of future VFI approach. In contrast to the kernel-based approach in Sec. II-B1, which encodes motion implicitly in spatially adaptive kernels and typically focuses on single-step CTFI, flow-based methods maintain an explicit dense flow field that can be reused to synthesize multiple timestamps and to inspect failure cases. This explicit representation provides strong interpretability and a large effective motion range, whereas kernel-based models, by tying their sampling patterns to local content, are often more robust in low-texture or motion-ambiguous regions but are less straightforward to extend to arbitrary timestamps or to diagnose at the level of explicit motion fields.
II-B3 Kernel- and Flow-based Combined